### v1

@@ Abstract
Stellar-mass black holes ($3$ $M_{\odot}\lesssim M_{\rm BH}\lesssim 150$ $M_{\odot}$) are the natural product of the evolution of heavy stars ($M_{\rm star}\gtrsim 20$ $M_{\odot}$). In our Galaxy, we expect $10^{8}$-$10^{9}$ stellar-mass black holes formed from the gravitational collapse of heavy stars, but currently we know less than 100 objects. We also know $\sim 100$ stellar-mass black holes in other galaxies, most of them discovered by gravitational wave observatories in the past 10 years. The detection of black holes is indeed extremely challenging and possible only in very special cases. This article is a short review on the physics and astrophysics of stellar-mass black holes.
@@ 1 Introduction
    Supermassive black holes have masses in the range $10^{5}$-$10^{10}$ $M_{\odot}$ and are found at the center of every middle-size and large galaxy ($M_{\rm galaxy}>10^{11}$ $M_{\odot}$) [^1] Exceptions may be possible: the galaxy A2261-BCG is one of the largest galaxies known but does not seem to host any supermassive black hole at its center Postman:2012qq . Kormendy:1995er . In the case of small galaxies ($M_{\rm galaxy}<10^{10}$ $M_{\odot}$), the situation is more controversial: some small galaxies host a supermassive black hole at their center while other small galaxies do not seem to have any supermassive black hole Ferrarese:2006fd ; Gallo:2007xq . While heavy objects can naturally migrate to the center of a multi-body system, which explains why supermassive black holes are at the center of their host galaxies, we do not know exactly how they formed and evolved: for example, they could have formed as stellar-mass black holes from the collapse of heavy stars and grown later or they could have formed from the collapse of heavy clouds and been heavier than stellar-mass black holes from the very beginning.
@@
where the sign -- ($+$) refers to co-rotating (counter-rotating) orbits and $Z_{1}$ and $Z_{2}$ are defined as
@@ 2 Stellar-Mass Black Holes
Stellar-mass black holes can be expected as the natural product of the evolution of heavy stars and their existence is confirmed by observations. When a star exhausts all its nuclear fuel, it cools down, the thermal pressure of its particles cannot compensate any longer the star’s own weight, and the body collapses. This is normally a violent process, which naturally ejects a significant fraction of material to space. As the star collapses, its density increases, and at some point electrons become degenerate: if the quantum pressure of electrons can stop the collapse, we have the formation of a white dwarf (this is the case of stars with $M_{\rm star}\lesssim 8$ $M_{\odot}$). The maximum mass for a white dwarfs is $\sim 1.4$ $M_{\odot}$ (Chandrasekhar limitChandrasekhar:1931ih . If the collapsing part of the star exceeds the Chandrasekhar limit, the quantum pressure of electrons cannot stop the collapse, the density increases further, and it is energetically convenient to convert protons and electrons into neutrons. At sufficiently high densities, neutrons become degenerate: if the quantum pressure of neutrons can stop the collapse, we have the formation of a neutron star (this is the case of stars with $8$ $M_{\odot}\lesssim M_{\rm star}\lesssim 20$ $M_{\odot}$). The maximum mass for a neuron star is $\sim 3$ $M_{\odot}$ (Oppenheimer-Volkoff limitOppenheimer:1939ne ; Rhoades:1974fn ; Kalogera:1996ci ; Lattimer:2012nd . If the neutron star exceeds the Oppenheimer-Volkoff limit, there is no known mechanism capable of stopping the collapse and we have the formation of a black hole (the case of stars with $M_{\rm star}\gtrsim 20$ $M_{\odot}$). The value of the minimum mass of a stellar-mass black hole formed from the collapse of a star corresponds thus to the maximum mass of a neutron star.
@@
The value of the maximum mass of a stellar-mass black hole formed from the collapse of a star depends on the mass of the progenitor star and its metallicity Heger:2002by ; Spera:2015vkd ; Mapelli:2021taw . Since heavy elements have larger cross sections than light elements, a higher metallicity increases the mass loss during the explosion of the star. Heavy stars with Solar metallicity can produce black holes with masses up to $\sim 20$ $M_{\odot}$ and the rest of the mass is ejected to space. Heavy stars with very low metallicity can collapse and produce black hole remnants with roughly half of their masses. The formation of black holes with masses between $\sim 60$ and $\sim 120$ $M_{\odot}$ (from metal-poor stars with masses between $\sim 120$ and $\sim 250$ $M_{\odot}$) is suppressed by pair-instability Heger:2001cd ; Belczynski:2016jno ; Woosley:2016hmi ; Woosley:2019nnp , but there are uncertainties on this map gas Farmer:2019jed ; Mapelli:2019ipt ; Costa:2020xbc ; Vink:2020nak . Black holes with masses above $\sim 150$ $M_{\odot}$ are unlikely simply because stars with $M_{\rm star}\gtrsim 300$ $M_{\odot}$ would exceed their Eddington luminosity and therefore cannot form [^3] Black holes with masses exceeding $\sim 150$ $M_{\odot}$ can still form from the merger of two black holes LIGOScientific:2025rsn ..
@@
From population evolution studies, we expect that in our Galaxy there are around $10^{8}$-$10^{9}$ stellar-mass black holes formed from the gravitational collapse of heavy stars. For example, assuming that the maximum mass for neutron stars is 1.7 $M_{\odot}$, Timmes et al. (1996) Timmes:1995kp predicted $1.4\cdot 10^{9}$ stellar-mass black holes in the Galaxy. Today we know that the maximum mass for a neutron star is somewhat higher Antoniadis:2013pzd , but from the model in Timmes et al. (1996) we can still expect around $1.0\cdot 10^{9}$ stellar-mass black holes in the Galaxy. On the other hand, Olejak et al. (2020) Olejak:2019pln predicted around $1.1\cdot 10^{8}$ stellar-mass black holes in the Galaxy: they estimated that in the Galactic disk there are about $1.0\cdot 10^{8}$ isolated black holes and about $8\cdot 10^{6}$ black holes in binary systems, in the Galactic bulge there are about $1.7\cdot 10^{7}$ isolated black holes and about $1\cdot 10^{6}$ black holes in binary systems, and in the Galactic halo there are about $4\cdot 10^{6}$ isolated black holes and about $5\cdot 10^{5}$ black holes in binary systems [^4] Most stellar-mass black holes are expected to be isolated rather than in a binary system with companion stars for the combination of three effects: $i)$ about 30% of massive stars are isolated stars, without companions; $ii)$ in a close binary, the two stars may merge before the supernova explosion of one of them, and $iii)$ in a wide binary, the supernova explosion can easily break up the system, producing an isolated black hole.
@@
We also know a few stellar-mass black holes in X-ray binaries and astrometric binaries in nearby galaxies and about 100 gravitational wave events associated to the coalescence of two stellar-mass black holes or the coalescence of a stellar-mass black hole and a neutron star. These objects will also be discussed briefly in the next subsections.
@@ 2.1 Black Holes in X-ray Binary Systems
Black hole X-ray binaries can be grouped into two classes: low-mass X-ray binaries ($M_{\rm star}\lesssim M_{\odot}$) and high-mass X-ray binaries ($M_{\rm star}\gtrsim 3$ $M_{\odot}$), where $M_{\rm star}$ is the mass of the companion star, not that of the black hole. In the known black hole X-ray binaries, the masses of the black holes are between $\sim 5$ $M_{\odot}$ and $\sim 20$ $M_{\odot}$.
@@
Figure 3: Cumulative histogram of the number of discovered stellar-mass black holes in transient X-ray sources. Red bars are for new black holes and blue bars are for dynamically confirmed black holes. The horizontal gray bars show the main X-ray missions used to discover and study these black holes. Figure from the online BlackCAT catalog https://www.astro.puc.cl/BlackCAT/ of Ref. Corral-Santana:2015fud .
@@
Figure 4: Sketch of 22 X-ray binaries with a stellar-mass black hole confirmed by dynamical measurements. For every system, the black hole accretion disk is on the left and the companion star is on the right. The color of the companion star roughly indicates its surface temperature (from brown to white as the temperature increases). The orientation of the disks indicates the inclination angles of the binaries. For comparison, in the top left corner of the figure we see the system Sun-Mercury: the distance between the two bodies is about 50 millions km and the radius of the Sun is about 0.7 millions km. Figure courtesy of Jerome Orosz.
@@ 2.2 Black Holes in Astrometric Binary Systems
    GAIA BH1 El-Badry:2022zih ; Chakrabarti:2022eyq (mass $M_{\rm BH}=9.62\pm 0.18$ $M_{\odot}$; orbital period of the binary $P\sim 185.6$ days; distance of the source $D\sim 480$ pc). As of now, this is the closest known black hole.
@@
    GAIA BH2 Tanikawa:2022xel ; El-Badry:2023pah (mass $M_{\rm BH}=8.9\pm 0.3$ $M_{\odot}$; orbital period of the binary $P\sim 1277$ days; distance of the source $D\sim 1.16$ kpc).
@@
    GAIA BH3 Gaia:2024ggk (mass $M_{\rm BH}=32.7\pm 0.8$ $M_{\odot}$; orbital period of the binary $P\sim 11.6$ years; distance of the source $D\sim 590$ pc). As of now, this is the heaviest known stellar-mass black hole in our Galaxy and the black hole binary with the longest orbital period.
@@
The discovery of these four black hole astrometric binaries is probably only the beginning and we can expect the discovery of many more black holes in the next years Mashian:2017kyp ; Breivik:2017cmy ; Wiktorowicz:2020dvb ; Janssens:2023tyt .
@@ 2.3 Isolated Black Holes
The identification of isolated black holes is definitively more challenging. As of now, we know only a robust candidate, which is associated to the microlensing event MOA-2011-BLG-191/OGLE-2011-BLG-0462 OGLE:2022gdj . The mass and the distance of this black hole are $M_{\rm BH}=7.1\pm 1.3$ $M_{\odot}$ and $D=1.58\pm 0.18$ kpc, respectively.
Murchikova & Sahu (2025) Murchikova:2025oio have recently proposed to look for isolated black holes accreting from the interstellar medium. The mass accretion rate of an isolated black hole moving through the interstellar medium can be estimated as
@@
where $v_{\rm BH}$ is the velocity of the black hole with respect to the interstellar medium, $c_{\rm ISM}$ is the sound speed in the interstellar medium, $\mu_{\rm ISM}$ is the mean atomic mass of the interstellar medium, $n_{\rm ISM}$ is the particle number density of the interstellar medium, and $m_{p}$ is the proton mass. Here $\lambda$ is a phenomenological parameter to describe the efficiency of the accretion process: $\lambda=1$ corresponds to the ideal Bondi accretion but its value is normally lower because of the presence of outflows (we can expect $10.01<\lambda<$ 1). Since $v_{\rm BH}\sim 50$ km s-1, the possibility of detecting isolated black holes accreting from the interstellar medium mainly depends on their distance from us, the value of $n_{\rm ISM}$, and sometimes even on the temperature of the interstellar medium through $c_{\rm ISM}$. The main challenge is to correctly identify these sources as accreting black holes. Indeed many isolated black holes may already be present in existing catalogs but have not yet been identified as such Murchikova:2025oio .
@@ 2.4 Black Holes in Compact Binary Systems
Black holes in binary systems black hole-black hole or black hole-neutron star can be detected by current gravitational wave observatories just before the merger of the two components, when the gravitational wave signal enters the sensitivity band of our current detectors (between a few Hz and a few kHz) and is sufficiently strong. Since the first detection in September 2015 LIGOScientific:2016aoc , about 100 events have been detected and the data are already public KAGRA:2021vkt . All these events were detected in the Observing Runs O1, O2, and O3 (see Fig. 5). The results of the Observing Run O4 will be released soon in September 2025 and should include over 200 events, roughly corresponding to the detection of a new event every 3 days.
@@ 3.1 Mass Gap
From the measurements of masses of neutron stars and black holes in X-ray binaries, we find that compact objects with masses between $\sim 2$ $M_{\odot}$ and $\sim 5$ $M_{\odot}$ are rare (see even Fig. 2Bailyn:1997xt ; Ozel:2010su ; Farr:2010tu . This would be a gap in the mass spectrum of compact objects formed from gravitational collapse, between the heaviest neutron stars and the lightest black holes. It is currently unclear if this mass gap is real or not. If it is real, it should be caused by the exact supernova explosion mechanism and may suggest a rapidly evolving explosion through a Rayleigh-Taylor instability Fryer:2011cx ; Belczynski:2011bn ; Fryer:2022lla . Otherwise, the gap may not be real and be the result of, for example, selection effects Siegel:2022gwc or systematic errors in current mass measurements Kreidberg:2012ud . The existence of such a mass gap has also been explored with gravitational wave observations Farah:2021qom ; KAGRA:2021duu , where compact objects in the mass gap have been reported LIGOScientific:2024elc . From the available gravitational wave events, we see a sharp peak around 8-10 $M_{\odot}$ in the black hole mass spectrum but no evidence of a true mass gap at 2-5 $M_{\odot}$ Ray:2025aqr .

### v2

@@ Abstract
Stellar-mass black holes ($3$ $_{\odot}\lesssim M_{\rm BH}\lesssim 150 MM_{\odot}$) are the natural product of the evolution of heavy stars ($_{\rm star}\gtrsim 20 MM_{\odot}$). In our Galaxy, we expect $10^{8}$-$10^{9}$ stellar-mass black holes formed from the gravitational collapse of heavy stars, but currently we know fewer than 100 objects. We also know $\sim 100$ stellar-mass black holes in other galaxies, most of them discovered by gravitational wave observatories in the past 10 years. The detection of black holes is indeed extremely challenging and possible only in very special cases. This article is a short review on the physics and astrophysics of stellar-mass black holes, including Galactic and extragalactic black holes in X-ray binaries, black holes in astrometric binaries, isolated black holes, and black holes in compact binaries. The article also addresses some important open issues and introduces the idea of a possible interstellar mission to the closest black hole.
@@ 1 Introduction
    Supermassive black holes have masses in the range $^{5}-1010$$^{10}$ $M_{\odot}$ and are found at the center of every middle-size and large galaxy ($_{\rm galaxy}>10^{11} MM_{\odot}$) [^1] Exceptions may be possible: the galaxy A2261-BCG is one of the largest galaxies known but does not seem to host any supermassive black hole at its center Postman:2012qq . Kormendy:1995er . In the case of small galaxies ($_{\rm galaxy}<10^{10} MM_{\odot}$), the situation is more controversial: some small galaxies host a supermassive black hole at their center while other small galaxies do not seem to have any supermassive black hole Ferrarese:2006fd ; Gallo:2007xq . While heavy objects can naturally migrate to the center of a multi-body system, which explains why supermassive black holes are at the center of their host galaxies, we do not know exactly how they formed and evolved: for example, they could have formed as stellar-mass black holes from the collapse of heavy stars and grown later or they could have formed from the collapse of heavy clouds and been heavier than stellar-mass black holes from the very beginning.
@@
where the sign - (++) refers to co-rotating (counter-rotating) orbits and $Z_{1}$ and $Z_{2}$ are defined as
@@ 2 Stellar-Mass Black Holes
Stellar-mass black holes can be expected as the natural product of the evolution of heavy stars Burrows:2023nlq ; Burrows:2024wqv and their existence is confirmed by observations. When a star exhausts all its nuclear fuel, it cools down, the thermal pressure of its particles cannot compensate any longer the star’s own weight, and the body collapses. This is normally a violent process, which naturally ejects a significant fraction of material to space.
As a star collapses, its density increases, and at some point electrons become degenerate: if the quantum pressure of electrons can stop the collapse, we have the formation of a white dwarf (this is the case of stars with $_{\rm star}\lesssim 8 MM_{\odot}$). The maximum mass for a white dwarfs is $\sim 1.4$ $M_{\odot}$ (Chandrasekhar limitChandrasekhar:1931ih . If the collapsing part of the star exceeds the Chandrasekhar limit, the quantum pressure of electrons cannot stop the collapse, the density increases further, and it is energetically convenient to convert protons and electrons into neutrons. At sufficiently high densities, neutrons become degenerate: if the quantum pressure of neutrons can stop the collapse, we have the formation of a neutron star (this is the case of stars with 88 $_{\odot}\lesssim M_{\rm star}\lesssim 20 MM_{\odot}$). The maximum mass for a neuron star is $\sim 3$ $M_{\odot}$ (Oppenheimer-Volkoff limitOppenheimer:1939ne ; Rhoades:1974fn ; Kalogera:1996ci ; Lattimer:2012nd . If the neutron star exceeds the Oppenheimer-Volkoff limit, there is no known mechanism capable of stopping the collapse and we have the formation of a black hole (the case of stars with $_{\rm star}\gtrsim 20 MM_{\odot}$). The value of the minimum mass of a stellar-mass black hole formed from the collapse of a star corresponds thus to the maximum mass of a neutron star.
@@
The value of the maximum mass of a stellar-mass black hole formed from the collapse of a star depends on the mass of the progenitor star and its metallicity Heger:2002by ; Spera:2015vkd ; Mapelli:2021taw . Since heavy elements have larger cross sections than light elements, a higher metallicity increases the mass loss during the explosion of the star. Heavy stars with Solar metallicity can produce black holes with masses up to $\sim 20$ $M_{\odot}$ and the rest of the mass is ejected to space. Heavy stars with very low metallicity can collapse and produce black hole remnants with roughly half of their masses. The formation of black holes with masses between $\sim 60$ and $\sim 120$ $M_{\odot}$ (from metal-poor stars with masses between $\sim 120$ and $\sim 250$ $M_{\odot}$) is suppressed by pair-instability Heger:2001cd ; Belczynski:2016jno ; Woosley:2016hmi ; Woosley:2019nnp , but there are uncertainties on this mass gap Farmer:2019jed ; Mapelli:2019ipt ; Costa:2020xbc ; Vink:2020nak . Black holes with masses above $\sim 150$ $M_{\odot}$ are unlikely simply because stars with $_{\rm star}\gtrsim 300 MM_{\odot}$ would exceed their Eddington luminosity and therefore cannot form [^3] Black holes with masses exceeding $\sim 150$ $M_{\odot}$ can still form from the merger of two black holes LIGOScientific:2025rsn ..
@@
From population evolution studies, we expect that in our Galaxy there are around $10^{8}$-$10^{9}$ stellar-mass black holes formed from the gravitational collapse of heavy stars. For example, assuming that the maximum mass for neutron stars is 1.7 $M_{\odot}$, Timmes et al. (1996) Timmes:1995kp predicted $1.4\cdot 10^{9}$ stellar-mass black holes in the Galaxy. Today we know that the maximum mass for a neutron star is somewhat higher Antoniadis:2013pzd , but from the model in Timmes et al. (1996) we can still expect around $1.0\cdot 10^{9}$ stellar-mass black holes in the Galaxy. On the other hand, Olejak et al. (2020) Olejak:2019pln predicted around $1.1\cdot 10^{8}$ stellar-mass black holes in the Galaxy: they estimated that in the Galactic disk there are about $1.0\cdot 10^{8}$ isolated black holes and about $8\cdot 10^{6}$ black holes in binary systems, in the Galactic bulge there are about $1.7\cdot 10^{7}$ isolated black holes and about $1\cdot 10^{6}$ black holes in binary systems, and in the Galactic halo there are about $4\cdot 10^{6}$ isolated black holes and about $5\cdot 10^{5}$ black holes in binary systems [^4] Most stellar-mass black holes are expected to be isolated rather than in a binary system with companion stars for the combination of three effects: $i)$ about 30% of massive stars are isolated stars, without companions; ii)ii) in a close binary, the two stars may merge before the supernova explosion of one of them, and $iii)$ in a wide binary, the supernova explosion can easily break up the system, producing an isolated black hole.
@@
We also know around 10 stellar-mass black holes in X-ray binaries and astrometric binaries in nearby galaxies (around 20 black holes if we include even weak candidates) and about 100 gravitational wave events associated to the coalescence of two stellar-mass black holes or the coalescence of a stellar-mass black hole and a neutron star. These objects will also be discussed briefly in the next subsections.
@@ 2.1 Black Holes in X-ray Binary Systems
Black hole X-ray binaries can be grouped into two classes: low-mass X-ray binaries ($M_{\rm star}\lesssim M_{\odot}$) and high-mass X-ray binaries ($_{\rm star}\gtrsim 3 MM_{\odot}$), where $M_{\rm star}$ is the mass of the companion star, not that of the black hole. In the known black hole X-ray binaries, the masses of the black holes are between $\sim 5$ $M_{\odot}$ and $\sim 20$ $M_{\odot}$.
@@
Figure 3: Cumulative number of discovered Galactic stellar-mass black holes in transient X-ray sources (red bars) and cumulative number of the dynamically confirmed stellar-mass black holes (blue bars). The horizontal gray bars show the periods of activity of X-ray missions relevant for the discovery and study of black hole X-ray binaries. Figure from the online BlackCAT catalog https://www.astro.puc.cl/BlackCAT/ of Ref. Corral-Santana:2015fud .
@@
Figure 4: Sketch of 22 X-ray binaries with a stellar-mass black hole confirmed by dynamical measurements. For every binary, we see the accretion disk around the black hole on the left and the companion star on the right. All accretion disks are shown at the inclination angle of the orbit of the binary. The colors of the companion stars indicate the surface temperatures of those stars (from brown to white as the temperature increases). The sizes of these binaries can be compared with the distance Sun-Mercury (about 50 millions km) reported in the top left corner of the cartoon. LMC X-1 and LMC X-3 are in the Large Magellanic Cloud, M33 X-7 is in the galaxy M33, and all other X-ray binaries are in our Galaxy. Figure courtesy of Jerome Orosz.
@@ 2.2 Black Holes in Astrometric Binary Systems
    GAIA BH1 El-Badry:2022zih ; Chakrabarti:2022eyq (mass $_{\rm BH}=9.62\pm 0.18 MM_{\odot}$; orbital period of the binary $P\sim 185.6$ days; distance of the source $D\sim 480$ pc). As of now, this is the closest known black hole.
@@
    GAIA BH2 Tanikawa:2022xel ; El-Badry:2023pah (mass $_{\rm BH}=8.9\pm 0.3 MM_{\odot}$; orbital period of the binary $P\sim 1277$ days; distance of the source $D\sim 1.16$ kpc).
@@
    GAIA BH3 Gaia:2024ggk (mass $_{\rm BH}=32.7\pm 0.8 MM_{\odot}$; orbital period of the binary $P\sim 11.6$ years; distance of the source $D\sim 590$ pc). As of now, this is the heaviest known stellar-mass black hole in our Galaxy and the black hole binary with the longest orbital period.
@@
The discovery of these four black hole astrometric binaries is probably only the beginning and we can expect the discovery of many more black holes in the next years Mashian:2017kyp ; Breivik:2017cmy ; Wiktorowicz:2020dvb ; Janssens:2023tyt ; Chawla:2023zhf  [^7] Early studies predicted that we could discover around $10^{5}$ black holes in binaries with Gaia Mashian:2017kyp , but they neglected binary evolution. Many binaries are disrupted during or before the formation of their black hole Breivik:2017cmy . More recent studies predict the possibility of discovering around $10^{3}$ black holes in binaries with Gaia Chawla:2023zhf ..
@@ 2.3 Isolated Black Holes
The identification of isolated black holes is definitively more challenging. As of now, we know only a robust candidate, which is associated to the microlensing event MOA-2011-BLG-191/OGLE-2011-BLG-0462 OGLE:2022gdj . The mass and the distance of this black hole are $_{\rm BH}=7.1\pm 1.3 MM_{\odot}$ and $D=1.58\pm 0.18$ kpc, respectively. Other potential isolated black hole candidates associated to microlensing events were reported in Refs. Bennett:2001vh ; Poindexter:2005hp ; Wyrzykowski20 .
Isolated black holes can accrete from the interstellar medium and produce a detectable flux of radiation Shvartsman ; Meszaros75 ; McDowell85 ; Campana93 ; Fujita:1997fh ; Tsuna:2018abi ; Kimura:2021ayq ; Murchikova:2025oio . The mass accretion rate of an isolated black hole moving through the interstellar medium can be estimated as Petrich89
@@
where $v_{\rm BH}$ is the velocity of the black hole with respect to the interstellar medium, $c_{\rm ISM}$ is the sound speed in the interstellar medium, $\mu_{\rm ISM}$ is the mean atomic mass of the interstellar medium, $n_{\rm ISM}$ is the particle number density of the interstellar medium, and $m_{p}$ is the proton mass. $\lambda$ is a dimensionless parameter <1<1 that reduces the ideal Bondi accretion rate because of the presence of outflows and convection (we can expect $0.01<\lambda<1$ Petrich89 ; Kaaz:2022dsa ; Galishnikova:2024nsk ; Kim:2024zjb ). Since $v_{\rm BH}\sim 50$ km s-1, the possibility of detecting isolated black holes accreting from the interstellar medium mainly depends on their distance from us, the value of $n_{\rm ISM}$, and sometimes even on the temperature of the interstellar medium through $c_{\rm ISM}$. The main challenge is to correctly identify these sources as accreting black holes. Indeed many isolated black holes may already be present in existing catalogs but have not yet been identified as such Murchikova:2025oio .
@@ 2.4 Black Holes in Compact Binary Systems
Black holes in binary systems black hole-black hole or black hole-neutron star can be detected by current gravitational wave observatories just before the merger of the two components, when the gravitational wave signal enters the sensitivity band of our current detectors (between a few Hz and a few kHz) and is sufficiently strong. Since the first detection in September 2015 LIGOScientific:2016aoc , about 100 events have been detected and the data are already public KAGRA:2021vkt . All these events were detected in the Observing Runs O1, O2, and O3 (see Fig. 5). The results of the Observing Run O4 will be released soon in Fall 2025 and should include over 200 events, roughly corresponding to the detection of a new event every 3 days.
@@ 3.1 Mass Gap
From the measurements of masses of neutron stars and black holes in X-ray binaries, we find that compact objects with masses between $\sim 2$ $M_{\odot}$ and $\sim 5$ $M_{\odot}$ are rare (see even Fig. 2Bailyn:1997xt ; Ozel:2010su ; Farr:2010tu . This would be a gap in the mass spectrum of compact objects formed from gravitational collapse, between the heaviest neutron stars and the lightest black holes. It is currently unclear if this mass gap is real or not.
If the mass gap is real, it may be caused by the exact supernova explosion mechanism and may suggest a rapidly evolving explosion through a Rayleigh-Taylor instability Fryer:2011cx ; Belczynski:2011bn ; Fryer:2022lla . Since the mass gap is found among neutron stars and black holes in binaries, it may be also caused by the supernova natal kick and be a peculiarity of neutron stars and black holes in binaries, not of the whole neutron star and black hole population; neutron stars Podsiadlowski:2003py and black holes heavier than 10 $M_{\odot}$ Mandel:2020qwb ; Burrows:2023ffl may receive only small natal kicks on average, while low-mass black holes may receive larger natal kicks, which may disrupt their binary system Fryer:1999ht . On the other hand, if the gap is not real, it may be the result of, for example, selection effects Siegel:2022gwc or systematic errors in current mass measurements Kreidberg:2012ud .
The existence of such a mass gap has also been explored with gravitational wave observations Farah:2021qom ; KAGRA:2021duu , where compact objects in the mass gap have been reported LIGOScientific:2024elc . From the available gravitational wave events, we see a sharp peak around 8-10 $M_{\odot}$ in the black hole mass spectrum but no evidence of a true mass gap at 2-5 $M_{\odot}$ Ray:2025aqr . A similar conclusion may be suggested by Gaia data Fishbach:2025bjh , where so far we have 21 neutron star candidates, a $\sim 3.6$ $M_{\odot}$ black hole candidate Song:2024tqr , and the three black holes discussed in Subsection 2.2: we have a first peak in the population of neutron stars at 1-2 $M_{\odot}$ and a second peak for $\sim 10$ $M_{\odot}$ black holes, but no true mass gap at 2-5 $M_{\odot}$ Fishbach:2025bjh . Future gravitational wave observations and the search for black holes in Gaia’s fourth data release will permit us to increase the statistics and to formulate more robust conclusions about the compact object population at 2-5 $M_{\odot}$.